Find a formula for the linear function described below.
The graph of  y=v(x) intersects the graph of u(x)=1+x^3 (x cubed)  at x=-2 and x=3.
v(x)=

Answer by stanbon(75887)

Find a formula for the linear function described below.
The graph of  y=v(x) intersects the graph of u(x)=1+x^3 (x cubed)  at x=-2 and x=3.
v(x)= ?
-----
Find the two points::
u(2) = 1+2^3 = 9
u(3) = 1+3^3 = 28
-----------------------
Linear form:: y = mx + b
Find "m" and "b"::
m = (28-9)/(3-2) = 19
----
Substitute and solve for "b"::
9 = 19*2 + b
b = -9
--------
v(x) = 19x - 9
----------------
Cheers,
Stan H.
----------------